8

The demand function of a firm is

Q = 90 − 1,5P,

where P and Q represent the price and quantity, respectively. At what price is revenue a maximum?

a.        15

b.        90

c.        30                                                             

d.        270


1
Expert's answer
2021-06-14T14:23:29-0400

The given demand function is

"Q=90-1.5P"

Revenue (R)=Q.P

"R=(90-1.5P).P"

For the maximum review,

"\frac{dR}{dP}=0"

"\Rightarrow 90-3P=0"

"\Rightarrow P= 30"

Hence, price should be equal to 30.